OpenStudy (anonymous):

limit of the funtion :- {[(2^n)+ 1][(7^n)+(10^n)]}^(1/n) as n tends to infinity

OpenStudy (isaiah.feynman):

Please use the equation editor to write the question so we can solve it.

OpenStudy (anonymous):

I am new in this site. I don't know what it is..

OpenStudy (isaiah.feynman):

\[\lim_{n \rightarrow \infty} [(2^{n}+1)(7^{n}+10^{n})]^{\frac{ 1 }{ n }}\]Is that the question?

OpenStudy (anonymous):

ya...you are right...Thanks for converting it so..

OpenStudy (isaiah.feynman):

I can't fully remember, but I know the entire function will approach a certain number as n approaches infinity.

OpenStudy (anonymous):

Do you require the answer?

OpenStudy (abb0t):

If I remember correctly, remember that \(\sf \color{}{\frac{1}{\infty}=0}\), yes? You can also foil the inside and see what you get: I believe you should get \(\infty\)

OpenStudy (anonymous):

No , I am sure the answer is not \[\infty \].

OpenStudy (dumbcow):

\[= \lim (1+2^{n})^{1/n} * \lim (7^{n} +10^{n})^{1/n}\] multiply by forms of 1 ... 2/2 and 10/10 \[2\lim (\frac{1+2^{n}}{2^{n}})^{1/n}*10 \lim (\frac{7^{n}+10^{n}}{10^{n}})^{1/n}\] = 2*10 = 20

OpenStudy (anonymous):

the second limit is not equal to 1

OpenStudy (perl):

did you try a ln

OpenStudy (perl):